The number of work-related injuries per month in a manufacturing plant is known to follow a Poisson distribution with a mean of 2.5 work-related injuries a month. What is the probability that in a given month no work-related injuries occur? What is the probability that atleast one work-related injury occurs


1
Expert's answer
2021-01-26T04:13:19-0500

Let "X=" the number of work-related injuries per month: "X\sim Po(\lambda)."


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

Given "\lambda=2.5"


"P(X=0)=\dfrac{e^{-2.5}\cdot2.5^0}{0!}=e^{-2.5}\approx0.082085"

"P(X\geq1)=1-P(X=0)=1-e^{-2.5}"

"\approx0.917915"